Hi, I wonder if you can help i can solve log equations with numbers in front but i am struggling with equations like this one:
ln(2x-1) + ln(2x+1) = 1.5
Would it be possible to show the process of calculating equation.
Thanks

Answer by ankor@dixie-net.com(12692)

ln(2x-1) + ln(2x+1) = 1.5
:
Find the antilog of both sides, on the right; (e^(1.5) = 4.481689
Remember adding logs means mult so you have:
(2x-1)*(2x+1) = 4.481689
FOIL
4x^2 - 1 = 4.481689
4x^2 = 4.481689 + 1
4x^2 = 5.481689
x^2 = 5.481689/4
x^2 = 1.37
x = SqRt(1.37)
x = 1.17
:
:
Check:
ln(2(1.17)-1) + ln(2(1.17)+1) =
ln(2.34-1) + ln(2.34+1)
ln(1.34) + ln(3.34) =
Find what the nat log is, should = 1.5
.292669 + 1.20597 = 1.498, not 1.5 because of rounding off
: